OpenStudy (anonymous):

@ttb123456789 haha I'm sorry, can you help with 1 more thing?

OpenStudy (anonymous):

i'm here

OpenStudy (anonymous):

big question!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

@ttb123456789

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think the unti rate is 24/10=2.4

OpenStudy (anonymous):

i think it means the rate or with what speed it goes but i'm not sure

OpenStudy (texaschic101):

unit rate is how much EACH crawfish costs You can make a proportion 10/24 = 1/x --- (10 lbs to 24 oz = 1 lb to x oz) cross multiply 10x = 24 x = 24/10 x = 2 2/5 seasonings or 2.4 so your unit rate is 2 2/5 seasonings for every 1 lb of crawfish

OpenStudy (anonymous):

Thank you guys! @texaschic101 @ttb123456789

OpenStudy (texaschic101):

oops...2 2/5 oz of seasoning for every lb of crawfish

OpenStudy (texaschic101):

glad to help

OpenStudy (anonymous):